# UASTHN: Uncertainty-Aware Deep Homography Estimation for UAV Satellite-Thermal Geo-localization

### v1

@@ Abstract
Geo-localization is an essential component of Unmanned Aerial Vehicle (UAV) navigation systems to ensure precise absolute self-localization in outdoor environments. To address the challenges of GPS signal interruptions or low illumination, Thermal Geo-localization (TG) employs aerial thermal imagery to align with reference satellite maps to accurately determine the UAV’s location. However, existing TG methods lack uncertainty measurement in their outputs, compromising system robustness in the presence of textureless or corrupted thermal images, self-similar or outdated satellite maps, geometric noises, or thermal images exceeding satellite maps. To overcome these limitations, this paper presents UASTHN, a novel approach for Uncertainty Estimation (UE) in Deep Homography Estimation (DHE) tasks for TG applications. Specifically, we introduce a novel Crop-based Test-Time Augmentation (CropTTA) strategy, which leverages the homography consensus of cropped image views to effectively measure data uncertainty. This approach is complemented by Deep Ensembles (DE) employed for model uncertainty, offering comparable performance with improved efficiency and seamless integration with any DHE model. Extensive experiments across multiple DHE models demonstrate the effectiveness and efficiency of CropTTA in TG applications. Analysis of detected failure cases underscores the improved reliability of CropTTA under challenging conditions. Finally, we demonstrate the capability of combining CropTTA and DE for a comprehensive assessment of both data and model uncertainty. Our research provides profound insights into the broader intersection of localization and uncertainty estimation. The code and data is publicly available.
@@ I Introduction
In this study, we introduce Uncertainty-Aware Satellite-Thermal Homography Network (UASTHN), a sample and consensus-based Uncertainty Estimation (UE) framework for Deep Homography Estimation (DHE) in satellite-thermal geo-localization, enhancing TG robustness.Our main contributions are as follows:
@@
In summary, our research enhances the reliability of UAV geo-localization by integrating uncertainty-aware localization with satellite and thermal data.
@@ II Related Works
UAV Thermal Geo-localization. Thermal Geo-localization (TG) involves using thermal cameras on UAVs in conjunction with satellite maps to extract the locations of thermal images and determine the UAV’s position. Existing approaches employ two primary strategies: global matching [8, 13] and local matching. This work focuses specifically on local matching methods.
@@ III-C Deep Ensembles (DE) and Merge Functions
To reduce the computational cost of extra samples, we propose an early stopping mechanism. It uses all samples for the first $k$ iterations for UE, then switches to a single sample solely for homography estimation, maintaining accuracy while significantly cutting computational overhead, as proven in our results.
@@ IV Experimental Setup
Dataset. We use the Boson-nighttime real-world thermal dataset [13, 14], which includes paired satellite-thermal images and unpaired satellite images. The thermal images were captured with a Boson Thermal Camera Module during nighttime flights (9:00 PM to 4:00 AM), covering landscapes such as deserts, farms, and roads over $33\leavevmode\nobreak\ $\mathrm{k}\mathrm{m}^{2}$$ for thermal and $216\leavevmode\nobreak\ $\mathrm{k}\mathrm{m}^{2}$$ for satellite images. The dataset uses Bing Satellite Maps, with thermal images ($W_{T}=512$) aligned to satellite images ($W_{S}=1536$). It contains $10$K training, $13$K validation, and $27$K test pairs, with test data from different regions and times. Additionally, there are $160$K unpaired satellite images for optional thermal synthesis, excluding validation and test regions to evaluate generalization.
Baselines. We use the following DHE baselines: DHN [16], IHN [17], and STHN [14]. DHN introduces DHE with four-corner displacement, while IHN adds an iterative approach with a correlation module. STHN uses TGM [13] to synthesize thermal images and a two-stage method for DHE in TG. We apply UE only for the first stage. Direct Modeling (DM) method [37] is our data uncertainty baseline.
@@
Implementation Details. We set the iteration numbers for the homography networks to $K_{1}=6$ for the first stage and $K_{2}=6$ for the second stage, with a resizing width $W_{R}$ of 256 and a decay factor $\gamma$ of 0.85. The model $F_{H}$ is initially trained without CropTTA, followed by fine-tuning with it. For two-stage methods, we apply bounding box augmentation [14], perturbing the first-stage results by 64 pixels. During the evaluation, the bounding box width $W_{B}$ is expanded by 64 pixels. All uncertainty estimation methods are evaluated with 5 samples. Homography networks are trained using the AdamW optimizer [38], with a linear decay scheduler and a maximum learning rate of $1\times 10^{-4}$.
@@ V-A Ablation Study
Sample Numbers. Fig. 3(d) illustrates the evaluation results for different sample numbers. The curves indicate that the performance of CropTTA and DE converges when the sample numbers exceed 4 and 3, respectively. This suggests the minimal sample numbers required for optimal performance.
@@ VI Conclusions
Future works will explore the development of an adaptive rejection mechanism and evaluate robustness across a more diverse dataset, including daytime and seasonal variations.

### v2

@@ Abstract
Geo-localization is an essential component of Unmanned Aerial Vehicle (UAV) navigation systems to ensure precise absolute self-localization in outdoor environments. To address the challenges of GPS signal interruptions or low illumination, Thermal Geo-localization (TG) employs aerial thermal imagery to align with reference satellite maps to accurately determine the UAV’s location. However, existing TG methods lack uncertainty measurement in their outputs, compromising system robustness in the presence of textureless or corrupted thermal images, self-similar or outdated satellite maps, geometric noises, or thermal images exceeding satellite maps. To overcome these limitations, this paper presents UASTHN, a novel approach for Uncertainty Estimation (UE) in Deep Homography Estimation (DHE) tasks for TG applications. Specifically, we introduce a novel Crop-based Test-Time Augmentation (CropTTA) strategy, which leverages the homography consensus of cropped image views to effectively measure data uncertainty. This approach is complemented by Deep Ensembles (DE) employed for model uncertainty, offering comparable performance with improved efficiency and seamless integration with any DHE model. Extensive experiments across multiple DHE models demonstrate the effectiveness and efficiency of CropTTA in TG applications. Analysis of detected failure cases underscores the improved reliability of CropTTA under challenging conditions. Finally, we demonstrate the capability of combining CropTTA and DE for a comprehensive assessment of both data and model uncertainty. Our research provides profound insights into the broader intersection of localization and uncertainty estimation. The code and models are publicly available.
@@ I Introduction
In this study, we introduce Uncertainty-Aware Satellite-Thermal Homography Network (UASTHN), a sample and consensus-based Uncertainty Estimation (UE) framework for Deep Homography Estimation (DHE) in satellite-thermal geo-localization.Our main contributions are as follows:
@@ II Related Works
UAV Thermal Geo-localization (TG). TG involves using thermal cameras on UAVs in conjunction with satellite maps to extract the locations of thermal images and determine the UAV’s position. Existing approaches employ two primary strategies: global matching [8, 13] and local matching. This work focuses specifically on local matching methods.
@@ III-C Deep Ensembles (DE) and Merge Functions
To reduce the computational cost of extra samples, we propose an early stopping mechanism. It uses all samples for the first $k$ iterations for UE, then switches to a single sample solely for homography estimation, maintaining accuracy while significantly cutting computational overhead.
@@ IV Experimental Setup
Dataset. We use the Boson-nighttime real-world thermal dataset [13, 14], which includes paired satellite-thermal images and unpaired satellite images. The 8-bit thermal images were captured with a Boson Thermal Camera during nighttime flights (9:00 PM to 4:00 AM), covering landscapes such as deserts, farms, and roads over $33\leavevmode\nobreak\ $\mathrm{k}\mathrm{m}^{2}$$ for thermal and $216\leavevmode\nobreak\ $\mathrm{k}\mathrm{m}^{2}$$ for satellite images. The dataset uses Bing Satellite Maps, with thermal images ($W_{T}=512$) aligned to satellite images ($W_{S}=1536$). It contains $10$K training, $13$K validation, and $27$K test pairs, with test data from different regions and times. Additionally, there are $160$K unpaired satellite images for optional thermal synthesis, excluding validation and test regions to evaluate generalization.
Baselines. We use the following DHE baselines: DHN [16], IHN [17], and STHN [14]. DHN introduces DHE with four-corner displacement, while IHN adds an iterative approach with a correlation module. STHN uses TGM [13] to synthesize thermal images and a two-stage method for DHE in TG. We apply UE only for the first stage. The Direct Modeling (DM) method [37] is our data uncertainty baseline.
@@
Implementation Details. We set the iteration numbers for the homography networks to $K_{1}=6$ for the first stage and $K_{2}=6$ for the second stage, with a resizing width $W_{R}$ of $256$ and a decay factor $\gamma$ of $0.85$. We first train $F_{H}$ for $100$k steps without CropTTA, then fine-tune it for an additional $200$K steps with CropTTA. For two-stage methods, we apply bounding box augmentation [14], perturbing the first-stage results by 64 pixels. During the evaluation, the bounding box width $W_{B}$ is expanded by 64 pixels. All uncertainty estimation methods are evaluated with 5 samples. $F_{H}$ is trained using the AdamW optimizer [38], with a linear decay scheduler and a maximum learning rate of $1\times 10^{-4}$.
@@ V-A Ablation Study
Sample Numbers. Fig. 3(d) illustrates the evaluation results for different sample numbers. The curves indicate that the performance of CropTTA and DE converge when the sample numbers exceed 4 and 3, respectively. This suggests the minimal sample numbers required for optimal performance.
@@ VI Conclusions
Future works will explore the development of an adaptive rejection mechanism and evaluate it on diverse datasets, including daytime, seasonal variations, and dynamic objects.